OpenStudy (anonymous):

I need to derive the loop equation for an RC Circuit Q/C-IR=0 to find the equation V=V0e^(-t/RC) can someone walk me step by step through it.

OpenStudy (anonymous):

For current leaving the capacitor you have \[\huge V=IR=\frac{Q}{C} = -\frac{dq}{dt}R\] separate variables AND take reciprocals \[\huge \frac{dq}{Q}=-\frac{dt}{RC}\] Now integrate \[\huge \ln|Q|+C = -\frac{t}{RC}\] \[\huge \ln|Q| = -\frac{t}{RC}+C\] raise both sides to e \[\huge Q=Q_{0}e^{-\frac{t}{RC}}\] Divide both sides by C to get; \[\huge \frac{Q}{C}= \frac{Q_0}{C}e^{-\frac{t}{RC}}\] Now remember that \[\huge V= \frac{Q}{C}, V_0 = \frac{Q_0}{C}\] so we have \[\huge V = V_0e^{-\frac{t}{RC}}\]